### 12 Mar 2026 v1

@@ Abstract
Figure 1: Given video sequences captured from multiple free cameras, our method can recover dense dynamic scenes consistently and estimate camera poses accurately. We illustrate our full pipeline (top) alongside reconstruction results on four additional sequences (bottom).
@@ 1 Introduction
Consider a scenario in which multiple observers simultaneously capture a dynamic scene from diverse viewpoints. Accurately reconstructing dynamic scenes from these observations is challenging due to temporal dynamics and limited view overlap. In this work, we study the problem of scene reconstruction from multiple free cameras: Given multiple video inputs, how can we recover a dense 3D model at any time step and estimate camera poses?
This problem is of significant practical importance. Multi-camera capture is increasingly common in robotics [12], sports broadcasting [45], and consumer devices where multiple phones or action cameras record a shared event. Applications such as augmented reality, dynamic Gaussian splatting [1, 42], and multi-view video analysis [23] all benefit from accurate, consistent 3D reconstruction across cameras. Yet existing dynamic scene reconstruction methods are limited to single-camera setups [46, 26, 24, 38, 16]. For instance, MegaSAM [24]—a robust monocular visual SLAM—exploits only temporal connections within each camera, leaving cross-camera consistency unexploited. While multi-camera SLAM methods exist [21, 35, 14, 44, 12], they either assume rigid camera rigs with known extrinsics or focus on static global map construction, failing to address dynamic scenes with freely moving cameras.
@@
To address these challenges, we introduce a two-stage optimization framework. In the first stage, we extend single-camera SLAM to multiple cameras by constructing a spatiotemporal connection graph that links frames across cameras based on view overlap, enabling joint bundle adjustment with consistent scale. To overcome the challenges of multi-camera initialization, we introduce a wide-baseline initialization strategy using a feed-forward reconstruction model. This approach provides a unified scale anchor and provides a reliable starting point for subsequent pose optimization. In the second stage, with coarse camera poses obtained, we refine per-frame depth and camera poses by optimizing dense inter- and intra-camera consistency using wide-baseline optical flow correspondences.
@@ 2 Related work
Classical visual SLAM methods [28, 10, 9, 17, 13] are primarily designed for single-view monocular camera configurations. Many works extend this setting to multi-camera configurations. One line of research employs calibrated camera rigs [12, 32, 20, 25], in which multiple cameras are rigidly mounted with known extrinsic parameters. Each camera provides complementary views of the environment, enabling wider field-of-view coverage and improved robustness in challenging scenarios. Another direction leverages multiple cameras in the context of multi-agent or collaborative SLAM [6, 36, 44, 8], where each camera (or agent) maintains an independent local mapping process. The individually built static sub-maps are later fused together. Our method differs from the above-mentioned approaches in two primary respects. First, unlike calibrated camera-rig systems, we allow cameras to move freely without requiring known inter-camera extrinsics. Second, in contrast to multi-agent SLAM frameworks, which typically assume static environments, our objective is to reconstruct dense and dynamic scenes.
Structure from motion (SfM) processes unordered image inputs, reconstructing the scene and estimating camera poses at the same time. Though accurate, classical SfM methods [30, 33] only achieve sparse representation and are brittle to dynamic objects in the view. On the other hand, recent feed-forward 3D reconstruction models [37, 40, 43], trained on large-scale datasets, exhibit substantially improved robustness under sparse input conditions compared with classical SfM pipelines. Nevertheless, such models are typically memory-intensive and encounter difficulties when applied to long video sequences. Though some methods [5] utilize such models in a chunk-wise way to handle long videos, their results are not as good as global predictions [7].
@@
If given camera poses, the simplest way is to use monocular depth prediction models [31, 3] to predict per-frame depth. However, the monocular depth model alone can often generate flickering and inconsistent results along a video [18]. Previous works solve the problem either via consistent video depth optimization [18, 27] or directly train video depth prediction [4, 19, 15]. Recently, many works reconstructing dynamic scenes rely on “3D reconstruction foundation" models [11, 46, 26, 39, 22, 41] to predict 3D scenes and camera poses at the same time. For example, Monst3R [46] fine-tunes the Dust3R model on dynamic videos and optimizes camera poses for long-sequence videos. CUT3R [39] uses a stateful recurrent model to process video inputs, achieving significant memory consumption reduction compared to global attention computation. However, existing feed-forward reconstruction models are either memory-intensive or produce low-quality reconstructions. In contrast, our method decouples camera pose estimation from depth optimization, enabling better reconstruction quality and scalability to long video sequences. The most relevant work to ours is [29], which reconstructs a mesh of dynamic objects in the view; however, it requires a fixed camera setup and prior extrinsic calibration. In contrast, our work can reconstruct scenes from freely moving cameras.
@@ 3.1 Overview
where $\mathbf{I}_{i}^{t}\in\mathbb{R}^{H\times W\times 3}$ denotes the image captured by the $i$-th camera at timestamp $t$, and $\mathbf{K}_{i}$ is the intrinsic matrix for the $i$th camera, the objective is to estimate the per-frame camera states
@@ 3.2 Spatio-temporal multi-camera tracking
We utilize the learned correspondence estimation model from MegaSAM [24]. Given two frames $\mathbf{I}_{i}$ and $\mathbf{I}_{j}$, MegaSAM learns to predict the dense optical flow $\mathbf{f}_{i\to j}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 2}$ and weights $\mathbf{w}_{ij}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ with lower resolution $(H^{\prime}\times W^{\prime})$ in an iterative way. Given a grid of pixel coordinates $\mathbf{u}_{i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 2}$ in frame $I_{i}$, we can predict the flow-warped correspondence of $\mathbf{u}_{i}$ on the image $I_{j}$ by $\mathbf{u}_{ij}^{\text{flow}}=\mathbf{u}_{i}+\mathbf{f}_{i\to j}$. With the current estimated state, we can also project $\mathbf{u}_{i}$ directly on the image $I_{j}$ and achieve the reprojection $\mathbf{u}_{ij}^{\text{reproj}}$:
@@
where $\mathbf{u}_{ij}^{\text{reproj}}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ is the reprojected correspondence of $\mathbf{u}_{i}$ on frame $I_{j}$; $\mathbf{T}_{ij}=\mathbf{T}_{j}^{-1}\circ\mathbf{T}_{i}$; $\mathbf{d}_{i}$ is the disparity map of frame $\mathbf{I}_{i}$. Given a group of connected frames $\mathbf{\Omega}$, parameters $\mathbf{T}$ and $\mathbf{d}$ are optimized by minimizing the weighted re-projection error:
@@
To extend the monocular tracking framework to multiple cameras, one of our key contributions is to introduce a spatio-temporal connection graph. Our frame connection graph consists of three parts. Temporal connection $\mathbf{\Omega}^{\rm{temp}}$: For each separate camera alone, since its frames are sequential in time, adjacent frames tend to have enough overlap for reliable correspondence estimation. In this case, we follow common practice in one single-camera setting, maintaining a temporal window to hold the latest keyframes for each camera, which means $\mathbf{\Omega}^{\rm{temp}}=\bigcup_{i}\mathbf{\Omega}_{i}^{\rm{temp}}$ for each $i$-th camera. Spatial connection $\mathbf{\Omega}^{\rm{spat}}$: In addition to the temporal intra-connection for each single camera, the spatial inter-connection is also necessary to maintain the scale consistency and improve accuracy. Given the $N$ frames from all cameras at timestamp $t$, we decide the inter-camera connection by evaluating the overlap across different camera frames: given two RGB frames $\mathbf{I}_{i}^{t}$ and $\mathbf{I}_{j}^{t}$, we project the grid of coordinates $\mathbf{u}_{i}$ to the frame $\mathbf{I}_{j}^{t}$ by Eq.˜1 to achieve the corresponding pixels $\mathbf{u}_{ij}^{t}$ on the frame $\mathbf{I}_{j}$’s image plane. If most of the projected pixels (more than 75%) are within the size of the image, then we make a spatial connection. This is helpful because even though there are dynamic objects in the scene, at the same timestamp, from different views, the scene is still consistent in this instant. Spatio-temporal connection $\mathbf{\Omega}^{\text{st}}$: In addition to the intra-camera temporal connection and inter-camera spatial connection at timestamp $t$, we also exploit the historical information across different cameras. At timestamp $t_{0}$, keyframe $\mathbf{I}_{i}^{t_{0}}$ will try to evaluate $\mathbf{u}_{ij}^{t_{0}}$ with all other cameras’ inactive keyframes $\{\mathbf{I}_{j}^{t}|j=[1,N],j\neq i,t=[1,T^{\prime}],T^{\prime}<t_{0}\}$. If there is enough overlap, we also make a connection between two frames. To prevent memory explosion and make connections effectively while running, we implement a connection balance strategy: we set a maximum number of edges in the active window and allocate inter-camera connections evenly if they exist. If newly detected connections plus the existing connections are more than the maximum number, we remove the oldest ones from the active window. A graphic demonstration of the spatio-temporal graph can be found in Fig.˜3.
@@
To address this issue, we adopt the robust feed-forward scene reconstruction model VGGT [37] to initialize our system. Specifically, during initialization, we select the first $N_{\text{init}}$ frames from each camera stream and input them into VGGT to obtain initial camera pose estimates along with per-frame depth predictions $\mathbf{D}_{\text{VGGT}}$. These predictions provide a coarse but globally consistent geometric initialization across multiple cameras.
To further enhance tracking robustness and to obtain dense depth priors, similar to MegaSAM, we additionally incorporate monocular depth estimates during tracking. Concretely, we employ a monocular depth estimation model, UniDepth [31], to predict per-frame monocular depth maps $\mathbf{D}_{\text{mono}}$. Since monocular depth is only defined up to an affine transformation, we align the predicted depths to the initialization scale by estimating a global scale $s\in\mathbb{R}$ and offset $o\in\mathbb{R}$ through optimizing
@@ 3.3 Multiple-view scene consistency refinement
Depth refinement from monocular video has been extensively studied in prior work [18, 27, 24]. In this section, we extend these techniques to the multi-camera setting, exploiting both temporal consistency within each camera and spatial consistency across different cameras. To improve optimization stability, our refinement procedure consists of three phases.
@@
To enable robust multi-view optimization, we compute dense optical flow correspondences that go beyond the low-resolution flow used during tracking. We construct an augmented connection graph $\mathbf{\Omega}_{\text{refine}}$ that includes: (1) the spatio-temporal graph $\mathbf{\Omega}$ from tracking (Sec.˜3.2), and (2) additional dense temporal connections within each camera sequence. Specifically, motivated by prior monocular video depth optimization [18, 27, 24], for each frame $i$ in a camera’s sequence, we connect it to frames at offsets $\{+2,+4,+8\}$ when they exist, enabling longer-range temporal consistency constraints.
For all frame pairs in $\mathbf{\Omega}_{\text{refine}}$, we estimate dense optical flow using the wide-baseline flow estimation model UFM [47], which provides higher quality correspondences than the tracking-stage flow, particularly for large inter-frame motion and inter-camera baselines.
@@
$\mathbf{u}_{ij}^{\text{reproj}}=\mathbf{K}_{j}(\mathbf{T}_{ij}\circ\mathbf{K}_{i}^{-1}(\mathbf{u}_{i},s_{i}\mathbf{D}_{i}^{\text{mono}}+\beta_{i})),$ (7)
where $s_{i}\in\mathbb{R}$ and $\beta_{i}\in\mathbb{R}$ are per-frame scale and shift parameters optimized in Phase 1, and $\mathbf{D}_{i}^{\text{mono}}$ is the raw monocular depth prediction. Note that these per-frame parameters are independent of the global initialization scale $(s,o)$ from Eq.˜3. The disparity consistency term $\mathcal{L}_{\text{disp}}$ penalizes inconsistencies between forward-projected and predicted depths, helping to reduce depth flickering [27]:
$\mathcal{L}_{\text{disp}}=\frac{\sum_{(i,j)}(\left|\frac{\max(\mathbf{D}_{j}^{\text{flow}},\mathbf{D}_{j}^{\text{geom}})}{\min(\mathbf{D}_{j}^{\text{flow}},\mathbf{D}_{j}^{\text{geom}})}-1\right|\cdot c_{i}+\lambda\log c_{i})\cdot m_{ij}}{\sum_{(i,j)}m_{ij}}$ (8)
where $\mathbf{D}_{j}^{\text{flow}}$ is the depth warped from frame $i$ to frame $j$ via optical flow, while $\mathbf{D}_{j}^{\text{geom}}$ is the monocular depth prediction at frame $j$.
@@
Phase 1 establishes frame-level scale consistency but does not correct per-pixel depth errors in the monocular predictions. In Phase 2, we fix per-frame affine parameters $(s_{i},\beta_{i})$, and directly optimize the per-pixel depth values $\mathbf{D}_{i}$ and refine camera poses by minimizing the same reprojection loss $\mathcal{L}_{\text{reproj}}$. Since jointly optimizing camera poses and depths suffers from instability [24], we propose to optimize poses and depths in an alternating iterative manner. For camera pose optimization, we augment the reprojection loss with pose regularization terms to maintain temporal smoothness:

### 14 Mar 2026 v2

@@ Abstract
Figure 1: Given video sequences captured from multiple free cameras, our method can recover dense dynamic scenes consistently and estimate camera poses accurately. We illustrate our full pipeline (top row) alongside reconstruction results on four additional sequences (bottom row).
@@ 1 Introduction
Consider a scenario in which multiple observers simultaneously capture a dynamic scene from diverse viewpoints. Accurately reconstructing dynamic scenes from these observations is challenging due to temporal dynamics and limited view overlap. In this work, we study the problem of dynamic scene reconstruction from multiple free cameras: Given multiple video inputs, how can we recover a dense 3D model at any time step and estimate camera poses?
This problem is of significant practical importance. Multi-camera capture is increasingly common in robotics [heng2019project], sports broadcasting [yus2015multicamba], and consumer devices where multiple phones or action cameras record a shared event. Applications such as augmented reality, dynamic Gaussian splatting [azzarelli2025splatography, wang2025monofusion], and multi-view video analysis [li2025multi] all benefit from accurate, consistent 3D reconstruction across cameras. Yet existing dynamic scene reconstruction methods are limited to single-camera setups [zhang2024monst3r, lu2025align3r, li2025megasam, wang2025shape, huang2025vipe]. For instance, MegaSAM [li2025megasam]—a robust monocular visual SLAM—exploits only temporal connections within each camera, leaving cross-camera consistency unexploited. While multi-camera SLAM methods exist [lajoie2020door, tian2022kimera, hu2023cp, yugay2025magic, heng2019project], they either assume rigid camera rigs with known extrinsics or focus on static global map construction, failing to address dynamic scenes with freely moving cameras.
@@
To address these challenges, we introduce a two-stage optimization framework. In the first stage, we extend single-camera SLAM to multiple cameras by constructing a spatio-temporal connection graph that links frames across cameras based on view overlap, enabling joint bundle adjustment with consistent scale. To overcome the challenges of multi-camera initialization under limited overlap, we introduce a wide-baseline initialization strategy using a feed-forward reconstruction model. This approach provides a unified scale anchor and provides a reliable starting point for subsequent pose optimization. In the second stage, with coarse camera poses obtained, we refine per-frame depth and camera poses by optimizing dense inter- and intra-camera consistency using wide-baseline optical flow correspondences.
@@ 2 Related work
Classical visual SLAM methods [mur2015orb, engel2014lsd, engel2017direct, klein2007parallel, homeyer2025droid] are primarily designed for single-view monocular camera configurations. Many works extend this setting to multi-camera configurations. One line of research employs calibrated camera rigs [heng2019project, schmied2023r3d3, kuo2020redesigning, liu2018towards], in which multiple cameras are rigidly mounted with known extrinsic parameters. Each camera provides complementary views of the environment, enabling wider field-of-view coverage and improved robustness in challenging scenarios. Another direction leverages multiple cameras in the context of multi-agent or collaborative SLAM [deng2025mne, 9686955, yugay2025magic, duhautbout2019distributed], where each camera (or agent) maintains an independent local mapping process. The individually built static sub-maps are later fused together. Our method differs from the above-mentioned approaches in two primary respects. First, unlike calibrated camera-rig systems, we allow cameras to move freely without requiring known inter-camera extrinsics. Second, in contrast to multi-agent SLAM frameworks, which typically assume static environments, our objective is to reconstruct dense and dynamic scenes.
Structure from motion (SfM) processes unordered image inputs, reconstructing the scene and estimating camera poses at the same time. Though accurate, classical SfM methods [pan2024global, schonberger2016structure] only achieve sparse representation and are brittle to dynamic objects in the view. On the other hand, recent feed-forward 3D reconstruction models [wang2025vggt, wang2024dust3r, yang2025fast3r], trained on large-scale datasets, exhibit substantially improved robustness under sparse input conditions compared with classical SfM pipelines. Nevertheless, such models are typically memory-intensive and encounter difficulties when applied to long video sequences. Though some methods [deng2025vggt] utilize such models in a chunk-wise way to handle long videos, their results are not as good as global predictions [ding2025laser].
@@
If given camera poses, the simplest way is to use monocular depth prediction models [piccinelli2024unidepth, bochkovskii2024depth] to predict per-frame depth. However, the monocular depth model alone can often generate flickering and inconsistent results along a video [kopf2021robust]. Previous works solve the problem either via consistent video depth optimization [kopf2021robust, luo2020consistent] or directly train video depth prediction [chen2025video, kuang2025buffer, hu2025depthcrafter]. Recently, many works reconstructing dynamic scenes rely on “3D reconstruction foundation" models [feng2025st4rtrack, zhang2024monst3r, lu2025align3r, wang2025continuous, leroy2024grounding, wang2025pi] to predict 3D scenes and camera poses at the same time. For example, Monst3R [zhang2024monst3r] fine-tunes the Dust3R model on dynamic videos and optimizes camera poses for long-sequence videos. CUT3R [wang2025continuous] uses a stateful recurrent model to process video inputs, achieving significant memory consumption reduction compared to global attention computation. However, existing feed-forward reconstruction models are either memory-intensive or produce low-quality reconstructions. In contrast, our method decouples camera pose estimation from depth optimization, enabling better reconstruction quality and scalability to long video sequences. The most relevant work to ours is [mustafa2016temporally], which reconstructs a mesh of dynamic objects in the view; however, it requires a fixed camera setup and prior extrinsic calibration. In contrast, our work can reconstruct scenes from freely moving cameras.
@@ 3.1 Overview
where $\mathbf{I}_{i}^{t}\in\mathbb{R}^{H\times W\times 3}$ denotes the image captured by the $i$th camera at timestamp $t$, and $\mathbf{K}_{i}$ is the intrinsic matrix for the $i$th camera, the objective is to estimate the per-frame camera states
@@ 3.2 Spatio-temporal multi-camera tracking
We utilize the learned correspondence estimation model from MegaSAM [li2025megasam]. Given two frames $\mathbf{I}_{i}$ and $\mathbf{I}_{j}$, MegaSAM learns to predict the dense optical flow $\mathbf{f}_{i\to j}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 2}$ and weights $\mathbf{w}_{ij}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ with lower resolution $(H^{\prime}\times W^{\prime})$ in an iterative way. Given a grid of pixel coordinates $\mathbf{u}_{i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 2}$ in frame $\mathbf{I}_{i}$, we can predict the flow-warped correspondence of $\mathbf{u}_{i}$ on the image $\mathbf{I}_{j}$ by $\mathbf{u}_{ij}^{\text{flow}}=\mathbf{u}_{i}+\mathbf{f}_{i\to j}$. With the current estimated state, we can also project $\mathbf{u}_{i}$ directly on the image $I_{j}$ and achieve the reprojection $\mathbf{u}_{ij}^{\text{reproj}}$:
@@
where $\mathbf{u}_{ij}^{\text{reproj}}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ is the reprojected correspondence of $\mathbf{u}_{i}$ on frame $\mathbf{I}_{j}$; $\mathbf{T}_{ij}=\mathbf{T}_{j}^{-1}\circ\mathbf{T}_{i}$; $\mathbf{d}_{i}$ is the disparity map of frame $\mathbf{I}_{i}$. Given a group of connected frames $\mathbf{\Omega}$, parameters $\mathbf{T}$ and $\mathbf{d}$ are optimized by minimizing the weighted re-projection error:
@@
To extend the monocular tracking framework to multiple cameras, one of our key contributions is to introduce a spatio-temporal connection graph. Our frame connection graph consists of three parts. Temporal connection $\mathbf{\Omega}^{\rm{temp}}$: For each separate camera alone, since its frames are sequential in time, adjacent frames tend to have enough overlap for reliable correspondence estimation. In this case, we follow common practice in one single-camera setting, maintaining a temporal window to hold the latest keyframes for each camera, which means $\mathbf{\Omega}^{\rm{temp}}=\bigcup_{i}\mathbf{\Omega}_{i}^{\rm{temp}}$ for each $i$-th camera. Spatial connection $\mathbf{\Omega}^{\rm{spat}}$: In addition to the temporal intra-connection for each single camera, the spatial inter-connection is also necessary to maintain the scale consistency and improve accuracy. Given the $N$ frames from all cameras at timestamp $t$, we decide the inter-camera connection by evaluating the overlap across different camera frames: given two RGB frames $\mathbf{I}_{i}^{t}$ and $\mathbf{I}_{j}^{t}$, we project the grid of coordinates $\mathbf{u}_{i}$ to the frame $\mathbf{I}_{j}^{t}$ by Eq.˜1 to achieve the corresponding pixels $\mathbf{u}_{ij}^{t}$ on the frame $\mathbf{I}_{j}$’s image plane. If most of the projected pixels (more than 75%) are within the size of the image, then we make a spatial connection. This is helpful because even though there are dynamic objects in the scene, at the same timestamp, from different views, the scene is still consistent in this instant. Spatio-temporal connection $\mathbf{\Omega}^{\text{st}}$: In addition to the intra-camera temporal connection and inter-camera spatial connection at timestamp $t$, we also exploit the historical information across different cameras. At timestamp $t_{0}$, keyframe $\mathbf{I}_{i}^{t_{0}}$ will try to evaluate $\mathbf{u}_{ij}^{t_{0}}$ with all other cameras’ inactive keyframes $\{\mathbf{I}_{j}^{t}|j=[1,N],j\neq i,t=[1,T^{\prime}],T^{\prime}<t_{0}\}$. If there is enough overlap, we also make a connection between two frames. To prevent memory explosion and make connections effectively while running, we implement a connection balance strategy: we set a maximum number of edges in the tracking active window and allocate inter-camera connections evenly if they exist. If newly detected connections plus the existing connections are more than the maximum number, we remove the oldest ones from the tracking active window. A graphic demonstration of the spatio-temporal graph can be found in Fig.˜3.
@@
To address this issue, we adopt the robust feed-forward scene reconstruction model VGGT [wang2025vggt] to initialize our system. Specifically, during initialization, we select the first $N_{\text{init}}$ frames from each camera stream and input them into VGGT to obtain initial camera pose estimates along with per-frame depth predictions $\mathbf{D}_{\text{VGGT}}$. These predictions provide a coarse but globally consistent geometric initialization across multiple cameras.
To further enhance tracking robustness and to obtain dense depth priors, similar to MegaSAM, we additionally incorporate monocular depth estimates during tracking. Concretely, we employ a monocular depth estimation model, UniDepth [piccinelli2024unidepth], to predict per-frame monocular depth maps $\mathbf{D}_{\text{mono}}$. Since monocular depth is only defined up to an affine transformation, we align the predicted depths to the initialization scale by estimating a global scale $s\in\mathbb{R}$ and offset $o\in\mathbb{R}$ through optimizing
@@ 3.3 Multiple-view scene consistency refinement
Depth refinement from monocular video has been extensively studied in prior work [kopf2021robust, luo2020consistent, li2025megasam]. In this section, we extend these techniques to the multi-camera setting, exploiting both temporal consistency within each camera and spatial consistency across different cameras. To improve optimization stability, our refinement procedure consists of two phases.
@@
To enable robust multi-view optimization, we compute dense optical flow correspondences that go beyond the low-resolution flow used during tracking. We construct an augmented connection graph $\mathbf{\Omega}_{\text{refine}}$ that includes: (1) the spatio-temporal graph $\mathbf{\Omega}$ from tracking (Sec.˜3.2), and (2) additional dense temporal connections within each camera sequence. Specifically, motivated by prior monocular video depth optimization [kopf2021robust, luo2020consistent, li2025megasam], for each frame $i$ in a camera’s sequence, we connect it to frames at offsets $\{+2,+4,+8\}$ when they exist, enabling longer-range temporal consistency constraints.
For all frame pairs in $\mathbf{\Omega}_{\text{refine}}$, we estimate dense optical flow using the wide-baseline flow estimation model UFM [zhang2025ufm], which provides higher quality correspondences than the tracking-stage flow, particularly for large inter-frame motion and inter-camera baselines.
@@
$\mathbf{u}_{ij}^{\text{reproj}}=\mathbf{K}_{j}(\mathbf{T}_{ij}\circ\mathbf{K}_{i}^{-1}(\mathbf{u}_{i},s_{i}\mathbf{D}_{i}+\beta_{i})),$ (7)
where $s_{i}\in\mathbb{R}$ and $\beta_{i}\in\mathbb{R}$ are per-frame scale and shift parameters to be optimized. Note that these per-frame parameters are independent of the global initialization scale $(s,o)$ from Eq.˜3. The disparity consistency term $\mathcal{L}_{\text{disp}}$ penalizes inconsistencies between forward-projected and estimated depths, helping to reduce depth flickering [luo2020consistent]:
$\mathcal{L}_{\text{disp}}=\frac{\sum_{(i,j)}(\left|\frac{\max(\mathbf{D}_{j}^{\text{flow}},\mathbf{D}_{j})}{\min(\mathbf{D}_{j}^{\text{flow}},\mathbf{D}_{j})}-1\right|\cdot\mathbf{c}_{i}+\lambda\log\frac{1}{\mathbf{c}_{i}})\cdot m_{ij}}{\sum_{(i,j)}m_{ij}}$ (8)
where $\mathbf{D}_{j}^{\text{flow}}$ is the depth warped from frame $i$ to frame $j$ via optical flow, while $\mathbf{D}_{j}$ is the estimated depth at frame $j$.
@@
Phase 1 establishes frame-level scale consistency but does not correct per-pixel depth errors in the monocular predictions. In Phase 2, we fix per-frame affine parameters $(s_{i},\beta_{i})$, and directly optimize the per-pixel depth values $\mathbf{D}_{i}$ and refine camera poses by minimizing the same reprojection loss $\mathcal{L}_{\text{reproj}}$. Since jointly optimizing camera poses and depths suffers from instability [li2025megasam], we propose to optimize poses and depths in an alternating iterative manner. For camera pose optimization, we augment the reprojection loss with pose regularization terms to maintain temporal smoothness: